# Rainbow chains and numerical renormalisation group for accurate chiral conformal spectra

### v1

@@ 3.2 Majorana fermions (critical Ising chain)
Figure 3: (a) Ground-state entanglement spectrum of the rainbow-deformed (3) TFI chain (8) of length $L=192$ and $\Delta=1/4$ in the even and odd fermion-parity sectors (blue and red), and after a boundary deformation to induce the SPT phase (green). The multiplicity and spacing of the degenerate subspaces correspond to the identity and fermion sectors of the Ising CFT in the trivial case, and a copy of the twist sector in each parity sector in the topological case.
(b) Entanglement spectrum of the rainbow-deformed (5) TFI ring for $L=192$ and $\Delta=1/4$. Colours indicate mirror-symmetry eigenvalues $+1$ (blue) and $-1$ (red) of the Schmidt vectors, which distinguish subspaces with even and odd momentum.
(c) Entanglement spectrum in the even-parity sector for a number of rainbow-deformed chains with fixed $\Delta=1/4$.
(d) Finite-size scaling of the $k=6,6.5,7,7.5$ multiplets on open chains as a function of the effective system size $w$ (cf. Appendix A). The deviation of each level from its expected value is proportional to $1/w^{2}$ (straight dotted lines as guides to the eye).
@@
The uniform ($J_{i}\equiv J$, $h_{i}\equiv h$) Hamiltonian has a transition between a ferromagnetic and a paramagnetic phase at $J=h$, which is described by the two-dimensional Ising CFT. In the entanglement spectrum of the corresponding rainbow-deformed chains, therefore, we expect to recover conformal towers of the chiral Ising CFT.
@@
$\gamma_{2i}$ $=c_{i}^{\dagger}+c_{i},$ $\gamma_{2i+1}$ $=i(c_{i}^{\dagger}-c_{i}).$ (9)
@@
In this Kitaev-chain [36] Hamiltonian, the TFI transition becomes one between a topologically trivial (for $h>J$) and a symmetry-protected topological (SPT, for $h<J$) phase, the latter of which has free Majorana edge modes. We found that wave functions belonging to the topological phase can be induced at the critical point by explicitly decoupling the first and last Majorana modes from the rest of the chain, allowing us to always work with the critical Hamiltonian in the bulk. In variational optimisation methods (e.g., DMRG), similar gossamer perturbations to drive the critical system in specific phases may be introduced by enforcing different on-site symmetries in the MPS [37].
@@
Entanglement spectra of TFI chains with couplings given by (3) are shown in Fig. 3(a) for boundary conditions consistent with both the trivial and the topological phase. In the former case, the entanglement spectrum consists of the conformal towers $(0)$ and $(1/2)$ of the $\mathcal{M}_{3}$ minimal model [13], which can be distinguished by the fermion parity of the corresponding Schmidt states [38]. In the topological phase, the entanglement spectrum in both parity sectors is identical and matches the $(1/16)$ conformal tower; the perfect twofold degeneracy of the entanglement spectrum, related to the presence of Majorana edge modes, is a hallmark of an SPT phase [2].
@@
The entanglement spectrum in the rainbow-deformed ring geometry (5) also recovers the $(0)\oplus(1/2)$ conformal towers expected for free boundary conditions [Fig. 3(b)]; as in the complex-fermion case, reflection eigenvalues of the corresponding Schmidt vectors correspond to the parity of the level of the conformal multiplets. Somewhat surprisingly, we were unable to find a deformation of this geometry under which the entanglement spectrum recovers the $(1/16)$ chiral conformal tower.
Similar to the tight-binding model in the previous section, we find that finite-size corrections to the predicted entanglement spectrum scale with the effective system size $w$ (Appendix A) as $1/w^{2}$, with only weak dependence on $\Delta$ [Figs. 3(c, d)].
@@ 4 Interacting systems: numerical renormalisation group
The MPO representing the Hamiltonian is folded at the strongest bond of the rainbow chain (a), so that sites with equal interaction strengths on the left and the right can be merged. The low-energy eigenstates of the Hamiltonian are likewise encoded in an MPS with doubled physical legs (b); using these, an environment tensor $L_{i}$, capturing the low-energy physics of the middle $2i$ sites of the chain, can be defined (c). The NRG step now consists of diagonalising the effective Hamiltonian (d); the lowest-energy $\chi_{\mathrm{NRG}}$ eigenvectors build up the new MPS tensor $M^{(i+1)}$, which is used to update the environment tensor (e), now describing a segment of length $2(i+1)$.
@@
Instead, we developed a numerical renormalisation-group (NRG) algorithm inspired by Wilson’s approach to the Kondo problem [23]. Here, the metallic bath was represented using a semi-infinite tight-binding chain (known as the Wilson chain) with exponentially decaying hopping terms, which allows for extracting the low-energy physics by iterative diagonalisation: Lowest-energy eigenstates of the first $(i+1)$ sites of the chain only have considerable overlap with the lowest-energy eigenstates of the chain of length $i$, since the relatively weak coupling terms to the last site do not introduce significant perturbations from higher energies. Therefore, the lowest-energy eigenstates at every chain length can be approximated by iteratively diagonalising and restricting the Hamiltonian in the next step onto the lowest few eigenstates. This process can naturally be written as the iterative construction of an MPS representation [Fig. 4(b)] of the lowest eigenstates [39, 40]: The action of the Hamiltonian [represented as a matrix-product operator (MPO)] on the lowest eigenstates of a chain of length $i$ can be compressed into an environment tensor [Fig. 4(c)], which yields the effective Hamiltonian for step $(i+1)$ [Fig. 4(d)]; the new MPS tensor is made up of the lowest-energy eigenvectors of this effective Hamiltonian, which (after scaling the MPO to counteract the exponential decay of Hamiltonian terms) also gives the new environment tensor for $(i+1)$ sites [Fig. 4(e)].
@@
This NRG protocol is popular for solving the Kondo problem and the Anderson impurity problem in dynamical mean-field theory [24], but has seen little use outside of this context. However, it is a convenient choice for solving rainbow chains, whose Hamiltonian terms also decay exponentially: The approximation of iterative diagonalisation and discarding high-energy states remains valid, even if the Hamiltonian is not a free-fermion one. Furthermore, as the rainbow chain represents a critical phase, we expect the NRG procedure to flow towards a fixed point, which can readily be verified numerically.
@@
The only change to the algorithm outlined above is that the Hamiltonian terms of equal strength on the two sides of the chain must be handled simultaneously, requiring us to fold the MPO Hamiltonian at the central bond of the rainbow chain [Fig. 4(a)]. [^2] This approach is identical to the standard NRG treatment of spinful baths, represented by two Wilson chains that only interact via the impurity site [23]. Accordingly, the low-energy states are represented by MPS whose tensors have two physical legs, each corresponding to a site in the left and right halves of the chain, respectively [Fig. 4(b)]. This in fact matches the entanglement structure of the expected rainbow ground state, which, in the limit $\alpha\to 0$, consists of singlet dimers precisely between the fused sites: such a state can be represented as an “MPS” of bond dimension 1. More generally, the folded representation naturally allows for volume-law entanglement, since the left and right halves of the chain are not coupled through a single MPS bond, but rather by all the tensors. This means that, in principle, we can obtain the ground state of a rainbow chain of arbitrary length and dynamical range of parameters, even outperforming the free-fermion method, which relies on numerically diagonalising the full single-particle Hamiltonian.
@@ 4.1 Example: the three-state Potts model
$E$ 2 $1\oplus 2$ +direct-sum+$\oplus$-+ ⊕ -
@@
In the absence of the boundary fields $h_{L,R}$, Eq. (10) is symmetric under simultaneous permutations of the basis states in (17) on all sites. These on-site symmetries form a non-Abelian $S_{3}$ group: eigenstates and entanglement spectra can be labelled with irreducible representations (irreps) of this group, which are listed in Table 1. In numerical simulations with TeNPy [41], we enforced the largest Abelian subgroup of $S_{3}$, the $\mathbb{Z}_{3}$ group of cyclic permutations. Irreps of this group are labelled by eigenvalues of $Z$, so the Hamiltonian can be written in the irrep basis by swapping the definitions of $X$ and $Z$ in (17). The boundary fields $h_{L,R}$, on the other hand, break the $\mathbb{Z}_{3}$ symmetry, leaving only the $\mathbb{Z}_{2}<S_{3}$ group generated by swapping the $\omega,\omega^{2}$ eigenvalues of $X$ (or, in fact, of $Z$). The correspondence between the irreps of $S_{3}$ and of these groups is also listed in Table 1.
The uniform (that is, $J_{i}\equiv J,h_{i}\equiv h$) Potts model has a self-dual critical point at $J=h$, described by the $\mathcal{M}_{5}$ conformal minimal model [42]. In analogy with the critical TFIM in Sec. 3.2, we therefore construct a rainbow chain with exponentially decaying couplings in the pattern shown in Fig. 1(c) (for easy comparison, we keep the parameter $\Delta$, defined through $\alpha=e^{-\Delta/2}$). We indeed find that for sufficiently high $\Delta$ and a sufficiently large number $\chi_{\mathrm{NRG}}$ of low-energy states kept, the NRG procedure converges to a fixed-point spectrum. This spectrum is shown in Fig. 5(a): In line with theoretical predictions for the three-state Potts CFT with free boundaries [16, 38], the spectrum consists of the conformal towers $(0)\oplus(3)$ (in the $0^{\pm}$ irreps of $S_{3}$) and two copies of the tower $(2/3)$ (in the two-dimensional $E$ irrep).
Figure 5: (a) Low-energy spectrum of the ferromagnetic critical 3-state Potts model on a rainbow chain, obtained from the numerical renormalisation group (NRG).
(b–c) Entanglement spectrum of the lowest-energy state of the rainbow chain in the ground state (b) and the first excited state in symmetry sector $E$ (c). Schmidt vectors are labelled with their eigenvalue under $\mathbb{Z}_{3}<S_{3}$, as well as their $\mathbb{Z}_{2}$ parity where that symmetry is preserved.
@@
(f–g) Scaling of the $k=6,6+2/3,7$ multiplets of the ground-state entanglement spectrum [grey rectangle in (b)] as a function of effective system size $w$ for a range of chain lengths $L$ (f) and of $\chi_{\mathrm{unzip}}$ (g). Finite-size effects scale as $1/\chi_{\mathrm{unzip}}^{2}$ and $1/w^{2}$ (dotted straight lines fitted to the data).
@@
Unless otherwise stated, all simulations use $L=256,\Delta=1/4,\chi_{\mathrm{NRG}}=\chi_{\mathrm{unzip}}=400$. Entanglement spectra are scaled by the first gap in the leading symmetry sector, and the lowest level is shifted to match the CFT prediction.
@@
We next obtained the entanglement spectrum of the ground state using the unzipping algorithm in Fig. 4(f–h) [Fig. 5(b)]: we find the same structure as in the energy spectrum, as expected. Since the ground state is symmetric under the full $S_{3}$ group, the irreps $0^{+}$ and $0^{-}$ can also be distinguished by computing the eigenvalues of the Schmidt vectors under the $\mathbb{Z}_{2}$ group generator, as shown in Fig. 4(i). We find that the $(0)$ and $(3)$ conformal towers appear in separate irreps: This is a remarkable difference to the partition function of the boundaryless three-state Potts model [43, 44], where these two conformal towers are bound together by offdiagonal terms.
@@
We also computed the entanglement spectrum of the lowest-energy state in the $\mathbb{Z}_{3}$ symmetry sector $Q=1$, shown in Fig. 5(c). This spectrum contains the conformal towers $(2/5)\oplus(7/5)$ ($Q=2$ sector) and two copies of the tower $(1/15)$ ($Q=0,1$ sectors): this is surprising, as these towers do not appear in the BCFT energy spectrum for free boundary conditions [38], but rather for certain fixed ones [16].
@@
In fact, entanglement spectra of states in nontrivial symmetry sectors do not map straightforwardly on a BCFT spectrum. Under the conformal transformation (2), the “ket” Hilbert spaces of the two halves of the rainbow-deformed chain in Fig. 1(b) map onto a “ket” Hilbert space at $\operatorname{Im}w=0$ and a “bra” Hilbert space at $\operatorname{Im}w=\pi$, respectively (cf. arrows in the figure). If these are related to one another by imaginary time evolution under a symmetric Hamiltonian, the symmetry quantum numbers of the “ket” Hilbert spaces must match; that is, the original chain must have trivial symmetry quantum numbers. Other entanglement spectra correspond to BCFT spectra with “twisted boundary conditions” in imaginary time, where symmetry quantum numbers change from $\operatorname{Im}w=0$ to $\pi$, then revert to their original values by $2\pi$. In future work, it will be interesting to understand the mechanism that gives rise to CFT spectra in such a setup, including predictions for the conformal towers that can be obtained this way.

### v2

@@ 3.2 Majorana fermions (critical Ising chain)
Figure 4: (a) Ground-state entanglement spectrum of the conformal (3) TFI chain (9) of length $L=192$ and $\Delta=1/4$ in the even and odd fermion-parity sectors (blue and red), and after a boundary deformation to induce the ferromagnetic/SPT phase (green). The number next to each CFT multiplet indicates its multiplicity; the numbers above each spectrum identify the corresponding conformal block of the Ising CFT: In the paramagnetic/trivial phase, we obtain the identity and fermion blocks; in the ferromagnetic/SPT one, two copies of the twist sector.
(b) Entanglement spectrum of the conformal (6) TFI ring for $L=192$ and $\Delta=1/4$. Colours indicate mirror-symmetry eigenvalues $+1$ (blue) and $-1$ (red) of the Schmidt vectors, which distinguish subspaces with even and odd momentum.
(c) Entanglement spectrum in the even-parity sector for several conformal chains with fixed $\Delta=1/4$.
(d) Schmidt values in the $h+n=6,6.5,7,7.5$ conformal levels [gray in panel (c)] as a function of effective system size $L_{\mathrm{eff}}$. Fits to the form $1/(aL_{\mathrm{eff}}^{2}+bL_{\mathrm{eff}}^{3})$ are added as guides to the eye.
@@
The uniform ($J_{i}\equiv J$, $h_{i}\equiv h$) Hamiltonian has a transition between a ferromagnetic and a paramagnetic phase at $J=h$, which is described by the two-dimensional Ising CFT. In the entanglement spectrum of the corresponding rainbow-deformed chains, therefore, we expect to recover conformal blocks of the chiral Ising CFT.
@@
$\gamma_{2i}$ $=c_{i}^{\dagger}+c_{i};$ $\gamma_{2i+1}$ $=i(c_{i}^{\dagger}-c_{i}),$ (10)
@@
as illustrated in Fig. 2. In this Kitaev-chain [37] Hamiltonian, the TFI transition becomes one between a topologically trivial (for $h>J$) and a symmetry-protected topological (SPT, for $h<J$) phase, the latter of which has free Majorana edge modes. We found that wave functions belonging to the topological phase can be induced at the critical point by explicitly decoupling the first and last Majorana modes from the rest of the chain, allowing us to always work with the critical Hamiltonian in the bulk. In variational optimisation methods (e.g., DMRG), similar gossamer perturbations to drive the critical system in specific phases may be introduced by enforcing different on-site symmetries in the MPS [38].
@@
Entanglement spectra of TFI chains with couplings given by (3) are shown in Fig. 4(a) for boundary conditions consistent with both the trivial and the topological phase. In the former case, the entanglement spectrum consists of the conformal blocks $(0)$ and $(1/2)$ of the $\mathcal{M}_{3}$ minimal model [13], which can be distinguished by the fermion parity of the corresponding Schmidt states [39]. In the topological phase, the entanglement spectrum in both parity sectors is identical and matches the $(1/16)$ conformal block; the perfect twofold degeneracy of the entanglement spectrum, related to the presence of Majorana edge modes, is a hallmark of an SPT phase [2].
@@
The entanglement spectrum in the rainbow-deformed ring geometry (6) also recovers the $(0)\oplus(1/2)$ conformal blocks expected for free boundary conditions [Fig. 4(b)]; as in the complex-fermion case, reflection eigenvalues of the corresponding Schmidt vectors correspond to the parity of the level of the conformal multiplets. Somewhat surprisingly, we were unable to find a deformation of this geometry under which the entanglement spectrum recovers the $(1/16)$ chiral conformal block.
3.3 Finite-size scaling
Figure 5: (a) Entanglement entropy $S_{\mathrm{vN}}$ of the tight-binding model (8) as a function of system size for five different values of $\Delta$ [indicated by colours, see inset and panel (b)]. The data for conformal rings is shifted by $+1$ to improve readability. They obey the volume law (11) (straight lines) to a good approximation. Inset: constant volume-law offset $\ell$ as a function of $\Delta$ for the five values of $\Delta$. They follow the expected scaling (11) (black lines).
(b) Entanglement gap $\delta_{\mathrm{ent}}$ in the dominant charge sector as a function of the effective system size $L_{\mathrm{eff}}=\frac{12}{c}S_{\mathrm{vN}}$. Data for the rainbow-chain and conformal ring geometries are shifted by ±$\pm 0.2$ to improve readability. For large $L_{\mathrm{eff}}$, the inverse proportionality (13) holds (solid lines); with a $1/L_{\mathrm{eff}}^{2}$ correction (dashed lines), it fits for every system size.
(c) Top: Deviation of single-particle entanglement energies $\varepsilon_{\alpha}$ ($\alpha=3/2$, $5/2$, $7/2$, $9/2$) from the thermodynamic-limit expectation for rainbow chains. Fits to the form $1/(aL_{\mathrm{eff}}^{2}+bL_{\mathrm{eff}}^{3})$ are added as guides to the eye. Bottom: Logarithmic derivative of the above for $\alpha=5/2$ and $9/2$.
(d–f) The same data for the transverse-field Ising model (9). Inset of (d): $\ell(\Delta)$ for the TFI chain matches $\ell(\Delta/2)$ for the tight-binding model (grey) very accurately.
As shown in Appendix A, the ground-state entanglement entropy of our models obeys a volume law with a $\Delta$-dependent correction:
$S_{\mathrm{vN}}$ $\simeq\frac{c}{12}[L\Delta+\ell(\Delta)];$ $\ell(\Delta)$$\simeq\begin{cases}-2\log\Delta+\mathrm{const.}&\text{chain geometries;}\\ -4\log\Delta+\mathrm{const.}&\text{ring geometries,}\end{cases}$ (11)
where $c$ is the central charge of the CFT, $c=1$ and $1/2$ for the tight-binding model (8) and the transverse-field Ising model (9), respectively. To check these predictions, we computed the entanglement spectrum for a range of system sizes ($8\leq L\Delta\leq 64$ with $\Delta=1/2$, $1/4$, $1/8$, $1/16$, $1/32$) for both Hamiltonians and all three geometries (3, 4, 6). Von Neumann entropies as a function of system size are plotted in Fig. 5(a, d), respectively: The prediction (11) is borne out perfectly in all cases. It is worth noting that $\ell(\Delta)$ for the TFIM matches $\ell(\Delta/2)$ for the tight-binding chain very closely: This underlines the notion that Hamiltonians with both bond and on-site terms are effectively discretised with doubled resolution (cf. Fig. 2). To control for the geometry- and $\Delta$-dependence of the entanglement entropy, we will use
$L_{\mathrm{eff}}=\frac{12}{c}S_{\mathrm{vN}}$ (12)
as our measure of system size from here on.
We also find analytically (Appendix B) that the entanglement gap $\delta_{\mathrm{ent}}$ within the leading charge sector is inversely proportional to the entanglement entropy for both Hamiltonians:
$\delta_{\mathrm{ent}}\simeq\frac{c\pi^{2}}{3S_{\mathrm{vN}}}\simeq\frac{4\pi^{2}}{L_{\mathrm{eff}}}.$ (13)
Our numerical results again match this prediction perfectly, albeit with stronger finite-size (but no visible finite-$\Delta$) effects [Fig. 5(b, e)].
We now move on to the structure of the entanglement spectrum. Since the ground states of both Hamiltonians are fermionic Gaussian states, their reduced density matrices can be written in the form
$\rho_{A}$ $=\prod_{\alpha}\left[\lambda_{\alpha}d_{\alpha}^{\dagger}d_{\alpha}+(1-\lambda_{\alpha})d_{\alpha}d_{\alpha}^{\dagger}\right].$ (14)
For tight-binding Hamiltonians, $\lambda_{\alpha}$ are the eigenvalues of the correlator matrix $\langle c_{i}^{\dagger}c_{j}\rangle$ restricted to subsystem $A$ [28], while mode creation operators follow from the eigenvectors $v_{\alpha}$ as $d_{\alpha}^{\dagger}=\sum_{i}(v_{\alpha})_{i}c_{i}^{\dagger}$. For Hamiltonians with pairing terms, we need to diagonalise the Nambu correlator matrix instead [30]. Due to Nambu symmetry, we get pairs of eigenvalues $(\lambda_{\alpha},\lambda_{\alpha^{\prime}}=1-\lambda_{\alpha})$ with mode operators $d^{\dagger}_{\alpha}=d_{\alpha^{\prime}}$: of course, only one from each pair should be included in (14).
In both cases, the entanglement spectrum and its scaling properties are governed by those of the “single-particle entanglement spectrum” $\lambda_{\alpha}$. It is useful to express these in terms of the the single-particle entanglement energies
$\varepsilon_{\alpha}=\log\frac{1-\lambda_{\alpha}}{\lambda_{\alpha}},$ (15)
which measure the change in entanglement energy upon filling the $d_{\alpha}^{\dagger}$ mode in a Schmidt state. Upon the conformal transformations (2, 5), these map onto the single-particle energies of an approximately uniform chain: Near the Fermi level, these have a sinusoidal dispersion that can be expanded as [20]
$\varepsilon_{\alpha}$ $\simeq A\alpha/(L\Delta)-B\alpha^{3}/(L\Delta)^{3}+\dots;$ $\alpha$ ={±$=\begin{cases}\pm\frac{1}{2},\pm\frac{3}{2},\dots&L=4n;\\ 0,\pm 1,\pm 2,\dots&L=4n+2.\end{cases}$ (16)
For large systems, higher-order terms of (16) are suppressed in powers of $1/L\Delta$, so the dispersion is well approximated as linear: In this limit, all Schmidt values within a conformal level coincide exactly. Numerically obtained entanglement energies indeed converge closer together with increasing $L$ [Fig. 3(e) and Fig. 4(c)], and $\varepsilon_{\alpha}/\delta_{\mathrm{ent}}$ converges to $\alpha$ [Fig. 5(c, f)].
[…]
$\frac{\varepsilon_{\alpha}}{\delta_{\mathrm{ent}}}-\alpha\simeq+L_{\mathrm{eff}}^{-3}.$ (17)
Accordingly, we expect that deviations of the entanglement spectrum from its expected thermodynamic-limit structure also scale as $L_{\mathrm{eff}}^{-3}$, since these are sums of single-particle entanglement energies $\varepsilon_{\alpha}$. However, the numerical results in Fig. 3(g) and 4(d) appear more consistent with the naïvely expected quadratic scaling. The reason for the discrepancy is that the cubic scaling only sets in for rather large systems [see the logarithmic derivatives in Fig. 5(c, f)], so for the system sizes that take up most of the figure, the quadratic scaling is a better approximation. In fact, fitting to the form $1/(aL_{\mathrm{eff}}^{2}+bL_{\mathrm{eff}}^{3})$ proved more accurate than either of the power laws.
@@ 4 Interacting systems: numerical renormalisation group
The MPO representing the Hamiltonian is folded at the strongest bond of the rainbow chain (a), so that sites with equal interaction strengths on the left and the right can be merged. The low-energy eigenstates of the Hamiltonian are likewise encoded in an MPS with doubled physical legs (b); using these, an environment tensor $L_{i}$ can be defined to capture the low-energy physics of the middle $2i$ sites of the chain (c). The NRG step now consists of diagonalising the effective Hamiltonian (d); the lowest-energy $\chi_{\mathrm{NRG}}$ eigenvectors build up the new MPS tensor $M^{(i+1)}$, which is used to update the environment tensor (e), now describing a segment of length $2(i+1)$.
@@
Instead, we developed a numerical renormalisation-group (NRG) algorithm inspired by Wilson’s approach to the Kondo problem [23]. Here, the metallic bath was represented using a semi-infinite tight-binding chain (known as the Wilson chain) with exponentially decaying hopping terms, which allows for extracting the low-energy physics by iterative diagonalisation: Lowest-energy eigenstates of the first $(i+1)$ sites of the chain only have considerable overlap with the lowest-energy eigenstates of the first $i$ sites, since the relatively weak coupling terms to the last site do not introduce significant perturbations from higher energies. Therefore, the lowest-energy eigenstates at every chain length can be approximated by iteratively diagonalising and restricting the Hamiltonian in the next step to the lowest few eigenstates.
This process can naturally be written as the iterative construction of an MPS representation [Fig. 6(b)] of the lowest eigenstates [40, 41]. The action of the Hamiltonian [represented as a matrix-product operator (MPO)] on the lowest eigenstates of a chain of length $i$ can be compressed into an environment tensor [Fig. 6(c)], which yields the effective Hamiltonian for step $(i+1)$ [Fig. 6(d)]. The new MPS tensor is made up of the lowest-energy eigenvectors of this effective Hamiltonian, which (after scaling the MPO to counteract the exponential decay of Hamiltonian terms) also gives the new environment tensor for $(i+1)$ sites [Fig. 6(e)].
@@
This NRG protocol is popular for solving the Kondo problem and the Anderson impurity problem in dynamical mean-field theory [24], but has seen little use outside of this context. Nevertheless, it is a convenient choice for solving rainbow chains, whose Hamiltonian terms also decay exponentially: The approximation of iterative diagonalisation and discarding high-energy states remains valid, even if the Hamiltonian is not a free-fermion one. Furthermore, as the rainbow chain represents a critical phase, we expect the NRG procedure to flow towards a fixed point, which can readily be verified numerically.
@@
The only change to the algorithm outlined above is that the Hamiltonian terms of equal strength on the two sides of the chain must be handled simultaneously, requiring us to fold the MPO Hamiltonian at the central bond of the rainbow chain [Fig. 6(a)]. [^1] This approach is identical to the standard NRG treatment of spinful baths, represented by two Wilson chains that only interact via the impurity site [23]. Accordingly, the low-energy states are represented by MPS whose tensors have two physical legs, corresponding to one site each in the left and right halves of the chain [Fig. 6(b)]. This in fact matches the entanglement structure of the expected rainbow ground state, which, in the limit $\alpha\to 0$, consists of singlet dimers precisely between the fused sites: such a state can be represented as an “MPS” of bond dimension 1. More generally, the folded representation naturally allows for volume-law entanglement, since the left and right halves of the chain are not coupled through a single MPS bond, but rather by all the tensors. This means that, in principle, we can obtain the ground state of a rainbow chain of arbitrary length and dynamical range of parameters, even outperforming the free-fermion method, which relies on numerically diagonalising the full single-particle Hamiltonian.
@@ 4.1 Example: the three-state Potts model
$E^{\phantom{-}}$ 2 $1\oplus 2$ +direct-sum+$\oplus$-+ ⊕ -
@@
In the absence of the boundary fields $h_{L,R}$, Eq. (18) is symmetric under simultaneous permutations of the basis states in (25) on all sites. These on-site symmetries form a non-Abelian $S_{3}$ group: eigenstates and entanglement spectra can be labelled with irreducible representations (irreps) of this group, which are listed in Table 1. In numerical simulations with TeNPy [42], we enforced the largest Abelian subgroup of $S_{3}$, the $\mathbb{Z}_{3}$ group of cyclic permutations. Irreps of this group are labelled by eigenvalues of $Z$, so the Hamiltonian can be written in the irrep basis by swapping the definitions of $X$ and $Z$ in (25). The boundary fields $h_{L,R}$, on the other hand, break this $\mathbb{Z}_{3}$ symmetry, leaving only the $\mathbb{Z}_{2}<S_{3}$ group generated by swapping the $\omega,\omega^{2}$ eigenvalues of $X$ (or, in fact, of $Z$). The correspondence between the irreps of $S_{3}$ and of these groups is also listed in Table 1.
The uniform (that is, $J_{i}\equiv J,h_{i}\equiv h$) Potts model has a self-dual critical point at $J=h$, described by the $\mathcal{M}_{5}$ conformal minimal model [43]. In analogy with the critical TFIM in Sec. 3.2, we therefore construct a rainbow chain with exponentially decaying couplings in the pattern shown in Fig. 1(c) (for easy comparison, we keep the parameter $\Delta$, defined through $\alpha=e^{-\Delta/2}$). We indeed find that for sufficiently high $\Delta$ and a sufficiently large number $\chi_{\mathrm{NRG}}$ of low-energy states kept, the NRG procedure converges to a fixed-point spectrum. This spectrum is shown in Fig. 7(a): In line with theoretical predictions for the three-state Potts CFT with free boundaries [16, 39], the spectrum consists of the conformal block $(0)\oplus(3)$ (in the $0^{\pm}$ irreps of $S_{3}$) and two copies of the block $(2/3)$ (in the two-dimensional $E$ irrep).
Figure 7: (a) Low-energy spectrum of the ferromagnetic critical 3-state Potts model on a rainbow chain (4), obtained from the numerical renormalisation group (NRG).
(b–c) Entanglement spectrum of the ground state (b) and the first excited state in symmetry sector $E$ (c), marked with filled symbols in panel (a). Schmidt vectors are labelled with their eigenvalue under $\mathbb{Z}_{3}<S_{3}$, as well as their $\mathbb{Z}_{2}$ parity where that symmetry is preserved.
@@
Unless otherwise stated, all simulations use $L=256$, $\Delta=1/4$, $\chi_{\mathrm{NRG}}=400$, $\chi_{\mathrm{unzip}}=800$. Entanglement spectra are scaled by the first gap in the leading symmetry sector, and the lowest level is shifted to match the scaling dimensions of the corresponding primary fields of $\mathcal{M}_{5}$.
@@
We next obtained the entanglement spectrum of the ground state using the unzipping algorithm in Fig. 6(f–h) [Fig. 7(b)]: we find the same structure as in the energy spectrum, as expected. Since the ground state is symmetric under the full $S_{3}$ group, the irreps $0^{+}$ and $0^{-}$ can also be distinguished by computing the eigenvalues of the Schmidt vectors under the $\mathbb{Z}_{2}$ group generator, as shown in Fig. 6(i). We find that the $(0)$ and $(3)$ conformal blocks appear in separate irreps: This is a remarkable difference to the partition function of the boundaryless three-state Potts model [44, 45], where these two blocks are bound together by offdiagonal terms.
@@
We also computed the entanglement spectrum of the lowest-energy state in the $\mathbb{Z}_{3}$ symmetry sector $Q=1$, shown in Fig. 7(c). This spectrum contains the conformal blocks $(2/5)\oplus(7/5)$ ($Q=2$ sector) and two copies of the block $(1/15)$ ($Q=0,1$ sectors): this is surprising, as these blocks do not appear in the BCFT energy spectrum for free boundary conditions [39], but rather for certain fixed ones [16].
@@
In fact, entanglement spectra of states in nontrivial symmetry sectors do not map straightforwardly on a BCFT spectrum. Under the conformal transformation (2), the “ket” Hilbert spaces of the two halves of the conformal chain in Fig. 1(b) map onto a “ket” Hilbert space at $\operatorname{Im}w=0$ and a “bra” Hilbert space at $\operatorname{Im}w=\pi$, respectively (cf. arrows in the figure). If these are related to one another by imaginary time evolution under a symmetric Hamiltonian, the symmetry quantum numbers of the “ket” Hilbert spaces must match; as a result, the original chain must have trivial symmetry quantum numbers. Other entanglement spectra correspond to BCFT spectra with “twisted boundary conditions” in imaginary time, where symmetry quantum numbers change from $\operatorname{Im}w=0$ to $\pi$, then revert to their original values by $2\pi$. In future work, it will be interesting to understand the mechanism that gives rise to CFT spectra in such a setup, including predictions for the conformal blocks that can be obtained this way.